OpenStudy (anonymous):

solve by the elimination method 5x+6y=3 10x+2y=6 and is this finite , infinate or no solution?

OpenStudy (anonymous):

5x+6y=3 = -10x-12y=-6 So cancel like terms -10x-12y=-6 10x+2y=6 add those together. -10y=0 so y=0 it is finite